OpenStudy (anonymous):

Jacob is having dinner at a restaurant. The menu includes 7 appetizers, 12 entrées, and 5 desserts. Jacob can select one appetizer, one entrée, and one dessert. What is the total number of dinner combinations that Jacob can select from?

OpenStudy (anonymous):

Wouldn't that be \[7C _{1}\times 12C _{1} \times 5C _{1}\] if you do it through the Combination method?

OpenStudy (anonymous):

Or, you could simply multiply \[7 \times 12 \times 5 \] to make things easier, since this doesn't seem like it requires us to use the formulae.